OpenStudy (sleepyjess):

How many grams of hydrochloric acid will react completely with a block of tin that is 2.5 cm by 3.0 cm by 3.5 cm, if the density of tin is 7.31 g/mL? Sn + HCl \(\rightarrow\) SnCl2 + H2

OpenStudy (sleepyjess):

@Kainui

OpenStudy (kainui):

Hey, alright so what ideas do you have so far to get started? I think the trickiest part is converting from cubic centimeters to mililiters. Try something out so I can help you where you're wrong. =)

OpenStudy (sleepyjess):

Isn't \(1cm^3\) the same as 1 mL?

OpenStudy (kainui):

Yep, I guess it was easier than I thought haha. The main idea is that tin and hydrochloric acid will react completely if there is an equal number of each.

OpenStudy (sleepyjess):

Sorry, I had to switch computers

OpenStudy (sleepyjess):

So it would need 7.31 g of hydrochloric acid?

OpenStudy (kainui):

Ok so what are we given and what do we want? We know we want grams of HCl, but we don't know how to get there quite yet. But we do know that there is a link from tin to HCl so let's see what we know about tin. So how much tin do we have in cubic centimeters?

OpenStudy (sleepyjess):

26.25

OpenStudy (abhisar):

Alright, so we have 26.25 \(\sf cm^3\) of tin. Now we need to convert this amount into ml using the density. can you do it?

OpenStudy (sleepyjess):

It would still be 26.25 mL I thought?

OpenStudy (abhisar):

oops! I mean you need to convert it into grams.

OpenStudy (abhisar):

Density is 7.31g/ml this means 1ml = 7.31g, so how much grams will be in 26.25 ml?

OpenStudy (sleepyjess):

26.25*7.31 191.8875

OpenStudy (abhisar):

\(\huge \checkmark\)

OpenStudy (sleepyjess):

Is that it?

OpenStudy (abhisar):

Now, balance the equation \(\sf Sn + HCl \rightarrow SnCl2 + H2\)

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @sleepyjess Is that it? \(\color{blue}{\text{End of Quote}}\) No this is only the amount of given tin. We need to find amount of HCl

OpenStudy (sleepyjess):

\(\sf Sn+2HCl\rightarrow SnCl_2+H_2\)

OpenStudy (abhisar):

\(\huge \checkmark\) Now, lets take a moment to understand what the question is asking. We have been given an equation and have been asked to find the amount of HCl that would be enough to completely react with the given tin block. The first step was to find the mass of the block in grams which we have successfully completed.

OpenStudy (abhisar):

ok?

OpenStudy (sleepyjess):

Ok

OpenStudy (sleepyjess):

The mass is 191.8875 right?

OpenStudy (abhisar):

Now lets take a look what information we can gather from the \(\bf balanced~equation\). \(\sf Sn+2HCl\rightarrow SnCl_2+H_2\) According to the equation 2 moles of HCl are required to react with 1 mole of tin. Getting it. \(\color{blue}{\text{Originally Posted by}}\) @sleepyjess The mass is 191.8875 right? \(\color{blue}{\text{End of Quote}}\) Yes.

OpenStudy (sleepyjess):

I think so

OpenStudy (abhisar):

Okay so what is the Atomic mass of tin ?

OpenStudy (abhisar):

@sleepyjess ?

OpenStudy (sleepyjess):

118.71

OpenStudy (abhisar):

Yes, so now you need to convert 191.8875 grams of tin into moles.

OpenStudy (sleepyjess):

\(\sf 191.8875*\dfrac{1~mol}{118.71}\)

OpenStudy (sleepyjess):

\(\sf 191.8875*\dfrac{1~mol}{118.71}\approx 1.6\)

OpenStudy (abhisar):

Great ! Now according to the equation, we saw that 2 moles of HCl is required to react with 1 mole of tin. Then how many moles of HCl are required to react with 1.6 moles of tin?

OpenStudy (sleepyjess):

1.6*2?

OpenStudy (abhisar):

Yes, now convert the answer into grams by using the molar mass of HCl.

OpenStudy (sleepyjess):

\(\sf 3.2*\dfrac{36.46}{1~mol}\approx 116.7\)

OpenStudy (abhisar):

\(\huge \checkmark\)

OpenStudy (sleepyjess):

:D

OpenStudy (sleepyjess):

\(\color{blue}{\text{Originally Posted by}}\) @sleepyjess Is that it? \(\color{blue}{\text{End of Quote}}\)

OpenStudy (sleepyjess):

Ok, now I am kinda confusing myself with this picture lol